A 18.0 kg child descends a slide 2.30 m high and reaches the bottom with a speed of 1.90 m/s . How much thermal energy due to friction was generated in this process?
Answer #1

The problem can be solved from energy method

Mass of child=18 kg

Height of slide=2.3 m

Potential energy of child with respect to bottom of slide=18*9.8*2.3=405.72 J

Speed of child at bottom of slide=1.9 m/s

Kinetic energy of child at bottom of slide

=0.5*18*1.9²=32.49 J

As we can see , the initial energy of child was 405.72 J amd final energy of child is 32.49 J

Therefore, the difference in energy is the enegy lost to friction

Thermal energy due to friction=405.72-32.49=373.23 J